A stone is dropped into a deep well and is heard to hit the water 2 s after being dropped. Determine the depth of the well. Assume no time delay in stone hitting the water and sound being heard. Take
a
=
10
m
s

2
.
1
Expert's answer
2021-08-02T08:40:21-0400

The distance travelled by the stone (the depth of the well) is given by the kinematic equation:


"h = \dfrac{at^2}{2}"

where "a =10m\/s^2" is the acceleration of the stone, and "t = 2s" is the time.

THus, obtian:


"h = \dfrac{10\cdot 2^2}{2} = 20m"

Answer. 20m.